# Conditional Generative Modeling for Enhanced Credit Risk Management in Supply Chain Finance

### v1

@@ 2 Functional Risk Formulation for Loan-Level Determination in SCF
Consider an SME engaged in cross-border e-commerce that sells products via an online platform and requires 3PL-led supply chain financing. In this setting, the goods are managed by a third-party logistics provider during transit, and the revenue generated from eventual sales is primarily allocated to repay the loan. As a financial institution partnering with the 3PL, our goal is to determine an appropriate loan amount $l$ based on both the collateral value and projected sales performance, ensuring adequate risk control while aligning financing with the SME’s revenue potential.
@@
where $(x)^{+}=\max\{x,0\}$. Intuitively, $L(l)$ captures the shortfall incurred when the loan amount $l$ exceeds the total revenue generated from sales, thereby encapsulating the risk associated with insufficient sales performance. This formulation abstracts the problem to its essential components, enabling us to evaluate how the loss (and hence the risk) varies as a function of the loan level.
The financial institution seeks to evaluate risk measures across the loan decision space $l\in[0,\bar{l}]$, where $\bar{l}$ denotes the maximum loan amount (often equal to the total collateral value to prevent arbitrage opportunities). We introduce a family of key risk measures that vary with $l$, offering flexibility in capturing different dimensions of credit risk. The first, the default probability, is defined as:
@@
representing the likelihood that sales revenue fails to cover the loan. The second, the expected loss, is given by:
@@
$\text{LGD}=\frac{\mathbb{E}[L(l)\mid L(l)>0]}{l}=\frac{\mathbb{E}[L(l)]}{l\,\Pr\{L(l)>0\}}=\frac{r_{2}(l)}{l\,r_{1}(l)}.$
Moreover, suppose we are interested in the probability that the loss exceeds a threshold level $\xi$, where $0<\xi<l$. Then,
$\Pr\{L(l)>\xi\}$ $=\Pr\{(l-rS)^{+}>\xi\}$
$=\Pr\{l-rS>\xi\}$
$=\Pr\{S<(l-\xi)/r\}$
$=r_{1}(l-\xi).$
and the expected loss above threshold $\xi$ is given by
$\mathbb{E}\left[L(l)\,\mathbb{I}\{L(l)>\xi\}\right]=$ $~{}\mathbb{E}\left[(l-rY)\,\mathbb{I}\{l-rY>\xi\}\right]$
$=$ $~{}\mathbb{E}\left[(l-\xi-rY)\,\mathbb{I}\{rY<l-\xi\}\right]$
@@
This demonstrate that $r_{1}(l)$ and $r_{2}(l)$ serve as building blocks for broader risk evaluations.
In addition, to further broaden the scope of our framework, we can also propose a more generalized risk measure, $r_{3}(l)$, defined as:
@@
where $g_{l}(\cdot)$ is a loan-level-dependent function that quantifies specific risk forms, and $a(l)$ is a threshold that varies with $l$, often tied to the default point (e.g., $a(l)=l/r$). This formulation further enhances flexibility, enabling $r_{3}(l)$ to adapt to a vastly wide range of risk metrics by tailoring $g_{l}(\cdot)$ and $a(l)$. The generality of $r_{3}(l)$ is most clearly evident in its ability to encompass both $r_{1}(l)$ and $r_{2}(l)$. By setting $g_{l}(Y)=1$ and $a(l)=l/r$, we obtain:
$r_{3}(l)=\mathbb{E}\left[1\cdot\mathbb{I}\{Y<l/r\}\right]=\Pr\{Y<l/r\}=r_{1}(l).$
Similarly, with $g_{l}(Y)=l-rY$ and $a(l)=l/r$, it follows that:
$r_{3}(l)=\mathbb{E}\left[(l-rY)\cdot\mathbb{I}\{Y<l/r\}\right]=\mathbb{E}[(l-rY)^{+}]=r_{2}(l).$
Beyond these foundational measures, $r_{3}(l)$ also extends to more complex risk assessments by allowing the loss function to take on more sophisticated, nonlinear forms. For instance, we can define $g_{l}(Y)=h(l-rY)$, where $h(x)=x^{p}$ for $p>0$ or other complex distortion functions. When $p=2$ and $a(l)=l/r$, this becomes $r_{3}(l)=\mathbb{E}[((l-rY)^{+})^{2}]$, which amplifies larger losses quadratically, emphasizing the severity of extreme shortfalls when sales fall well below the loan amount. Moreover, $g_{l}(Y)$ can be tailored to address scenario-dependent risks, such as those influenced by liquidity constraints or regulatory penalties by setting $g_{l}(Y)=(l-rY)\cdot(1+kY)$, where $k$ is an additional cost coefficient. Thus, through appropriate choices of $g_{l}(\cdot)$ and $a(l)$, $r_{3}(l)$ can offer a versatile framework for evaluating sophisticated risk measures, ensuring robustness and adaptability in diverse SCF applications and real-world financial challenges.
Now, the key challenge becomes efficiently computing these risk measures across all $l\in[0,\bar{l}]$. The crux of this challenge is to learn the distribution of $Y$. Suppose we have a generative model $\hat{Y}$ capable of producing $K$ independent and identically distributed observations $\{\hat{Y}_{i}\}_{i=1}^{K}$ rapidly, such that the distribution of $\hat{Y}$ closely approximates that of $Y$. Then, we can estimate the risk measures using Monte Carlo simulation (Hong et al. 2014):
@@
Our proposed functional risk assessment framework revolutionizes traditional SCF credit risk evaluations, which typically rely on methods like logistic regression or support vector machines to estimate a single risk measure—often the probability of default—for a fixed loan amount. By modeling risk measures such as $r_{1}(l)$, $r_{2}(l)$, and the versatile $r_{3}(l)$ as functions of loan size $l$, our functional approach provides a unified and comprehensive risk profile across a range of loan levels. Supported by a fast and reliable generative model and Monte Carlo approximations, this framework enables financial institutions to assess diverse credit risk measures and make informed loan sizing decisions, aligning financing with sales potential while significantly enhancing SCF risk management compared to static, single-point traditional methods.
@@ 3 Constructing a Conditional Generative Model for Sales Distribution
In this section, we first formulating the problem of constructing a conditional generative model for sale distribution. We then present QRGMM as an effective generative modeling technique that meets the specific requirements of SCF risk management. Finally, we establish theoretical guarantee for the learning of key risk measures, $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$, using QRGMM.
@@ 3.1 Problem Formulation
When a SME raises financing through 3PL-led CEBC-SCF mode, the sales revenue of its products will generally be prioritized to repay the loan. Hence, the random variable representing product sales, $Y(\mathbf{x})$, directly determines the loan repayment outcome. Here, the multi-dimensional vector $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{p}$ denotes the associated covariates, such as product features, seller characteristics, and other auxiliary information. We collect $n$ historical observations $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ from e-commerce platforms, where $y_{i}$ are realizations of $Y(\mathbf{x}_{i})$. To effectively manage credit risk in SCF, our objective is to develop a generative model $\hat{Y}(\mathbf{x})$ utilizing the dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. This model should accurately and robustly generate random observations that replicate the conditional distribution of $Y(\mathbf{x})$ given the covariates $\mathbf{x}$, while also ensuring rapid generation.
While generative models have significantly propelled advancements in areas such as image and video synthesis as well as language generation (Goodfellow et al. 2014, Radford et al. 2018, Ho et al. 2020, Liu et al. 2024), many of them are designed for high-dimensional data, focusing on visual or semantic realism rather than accurately capturing low-dimensional distributions, especially in the tails. Some models like GANs and variational autoencoders (Kingma and Welling 2013) also suffer from mode collapse issues (Saxena and Cao 2021), which reduces data diversity and fails to replicate the true distribution accurately, while others (e.g., diffusion-based approaches) are too slow for large-scale Monte Carlo simulation. In contrast, QRGMM (Hong et al. 2023) is designed to address these issues, offering a compelling alternative for SCF applications.
@@ 3.2 The QRGMM Algorithm
  • For $k=1,\ldots,K$, output $\hat{Y}_{k}(\mathbf{x})$ as follows:
    for $u_{k}\in[\tau_{j},\tau_{j+1})$ with $j=1,\ldots,m{-}2$,
@@
  • Accuracy and Robustness: By explicitly modeling conditional quantiles at multiple levels, QRGMM learns the full shape of a distribution rather than focusing merely on mean predictions or specific performance measures. This quantile-based approach inherently avoids issues such as mode collapse, and it provides more reliable estimates of the tails of the distribution which have more impact on the SCF credit risk. Unlike adversarial models such as GANs that require solving challenging saddle-point optimization (notoriously difficult to tune), QRGMM merely involves fitting quantile regressions and selecting the main parameter—the number of grid points $m$. Furthermore, the choice of $m=O(\sqrt{n})$, guided by theoretical and empirical insights of Hong et al. (2023), effectively balances performance with computational cost, resulting in a simpler and more stable training process.
@@
  • Speed: Numerical results in Hong et al. (2023) show that QRGMM can generate a lot of observations in an extremely short time—e.g., 100,000 random samples in under 0.01 seconds. This efficiency stems from its reliance on simple linear interpolation in the online generation stage, after the quantile regressions have been fitted offline. Such rapid sampling is particularly valuable in SCF risk management, as it enables real-time generation of extensive samples, significantly reducing the computational cost of sample generation. This allows us to produce large sample sets to support Monte Carlo methods, computing various risk measures of interest with enhanced flexibility, while minimizing estimation errors caused by sampling variance.
@@
In summary, by framing SCF risk management through the lens of the full conditional distribution of product sales, we underscore the need for a generative model that is not only accurate and robust but also computationally efficient. QRGMM effectively meets these requirements, providing a promising tool for the integrated framework that links credit risk evaluation with loan sizing decisions. However, to fully realize the practical potential of QRGMM in SCF applications, it is essential to go beyond its generative capabilities and rigorously establish how its learning outcomes translate into reliable credit risk measures. Specifically, building on the problem formulation in Section 2, the next critical step is to provide strong theoretical guarantees on the estimation of the key risk measures—namely, $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$—across the entire range of loan levels $l\in[0,\bar{l}]$. This ensures that the QRGMM-based framework is not only practically effective but also theoretically sound, setting the stage for robust real-world deployment.
3.3 Asymptotic Analysis of Risk Measure Estimations via QRGMM
@@
In this subsection, we establish theoretical guarantees for the estimation of $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$ for all $l\in[0,\bar{l}]$, thereby ensuring asymptotic validity for a broad range of SCF risk measure estimations through QRGMM. Prior to proceeding, we emphasize several key considerations.
@@
In this section, we adopt the linear quantile regression model assumption of Hong et al. (2023), which lay the theoretical groundwork for analyzing the asymptotic properties of QRGMM. Suppose $Y(\mathbf{x})$ is a continuous random variable on its support given the input $\mathbf{x}$. Its conditional distribution and density functions are denoted as $F_{Y}(\cdot|\mathbf{x})$ and $f_{Y}(\cdot|\mathbf{x})$. The linear quantile regression model assumption is stated as follows.
@@
$F^{-1}_{Y}(\tau|\mathbf{x})=\boldsymbol{\beta}(\tau)^{\intercal}\mathbf{x}$ all $\tau\in(0,1)$ and $\boldsymbol{\beta}(\tau)$ is continuous in $\tau$.
Assumption 1 posits that the linear quantile regression model accurately represents the true model of the conditional quantile functions of $Y$ given $\mathbf{x}$ for all $\tau$ in $(0,1)$, implying that $F^{-1}_{Y}(\tau|\mathbf{x})$ is linear with respect to the covariates $\mathbf{x}$. Although Assumption 1 may seem restrictive, it is widely used in practice due to its simplicity and interpretability. This assumption is foundational not only because it simplifies the theoretical framework but also because it is underpinned by a substantial literature on classical quantile regression, which provides the necessary theoretical foundation for establishing the asymptotic analysis. Furthermore, in conjunction with domain knowledge, it can be expanded to capture nonlinearity by substituting $\mathbf{x}$ with a set of nonlinear basis functions $\boldsymbol{b(\mathbf{x})}$. In terms of theoretical analysis, the theoretical results can be readily extended when employing basis functions. Supporting evidence from numerical experiments by Hong et al. (2023) demonstrates that even when the linear model assumption does not strictly hold, the approximation using basis functions still produces favorable outcomes. In Section 4, we will explore an extension of the linear quantile regression model to the DeepFM model, enhancing its ability to capture complex combinatory features.
@@
Source of Error
Our credit risk estimation procedure builds upon QRGMM to produce a generative model $\hat{Y}(\mathbf{x})$ that approximates the distribution of $Y(\mathbf{x})$. Given $\hat{Y}(\mathbf{x})$, we then rely on Monte Carlo simulation to estimate risk measures by generating $K$ samples $\{\hat{Y}_{k}(\mathbf{x})\}_{k=1}^{K}$ from the generative model and transforming them into the estimation of risk measures through (2). This process involves three distinct sources of approximation:
@@
    Interpolation Error: To approximate $F^{-1}_{Y}(\tau|\mathbf{x})$ for all $\tau$, we rely on a set of $m$ quantile levels and piecewise interpolation. This discretization, governed by $m$, introduces interpolation error that diminishes as we increase the number of quantile levels.
@@
    Monte Carlo Sampling Error: After constructing $\hat{Y}(\mathbf{x})$, the final estimates of risk measures are computed by sampling from this approximate distribution. The finite number $K$ of Monte Carlo samples leads to sampling variability, which decreases as $K$ grows.
@@
3.3.1 Asymptotic Analysis for $r_{1}(l)$ Estimation
In this subsection, we establish the asymptotic properties of the estimated default probability $\hat{r}_{1}(l)$, a key metric in credit risk assessment for SCF. Given any $\mathbf{x}$, the default probability $r_{1}(l)$ satisfies:
$r_{1}(l)=\Pr\{L(l)>0\}=\Pr\{Y(\mathbf{x})<l/r\}=F_{Y}\bigl{(}\,l/r\!\mid\!\mathbf{x}\bigr{)},$
In the QRGMM framework, ignoring the Monte Carlo sampling error, $r_{1}(l)$ is essentially approximated by $\hat{r}_{1}(l)=F_{\hat{Y}}(\,l/r|\mathbf{x})$, where $F_{\hat{Y}}(\cdot|\mathbf{x})$ is the CDF of the generative model $\hat{Y}(\mathbf{x})$. The accuracy of $\hat{r}_{1}(l)$ thus hinges on the convergence of $F_{\hat{Y}}(y|\mathbf{x})$ to the true CDF $F_{Y}(y|\mathbf{x})$.
In SCF context, by understanding the distribution of $Y(\mathbf{x})$, lenders can determine the appropriate loan level $l$ such that we can achieve a desired PD. To support this loan sizing decision making, it is not enough to have mere pointwise convergence of $F_{\hat{Y}}(\,l/r|\mathbf{x})$ for a fixed $l$. Instead, a stronger form of uniform convergence is required. Uniform convergence of $F_{\hat{Y}}(y|\mathbf{x})$ for all $y\in[0,\infty)$ ensures the uniform convergence of $\hat{r}_{1}(l)$ for all $l\in[0,\bar{l}]$. This uniform convergence is a stronger and more challenging requirement, as it ensures consistent accuracy across the entire loan decision space, addressing a gap in existing literature that often overlooks the responsive interplay between loan levels and risk measures.
Our objective is to prove both pointwise and uniform convergence of $\hat{r}_{1}(l)$ to $r_{1}(l)$. These theoretical guarantees are essential for ensuring the reliability of QRGMM in SCF. Leveraging results from Hong et al. (2023), we proceed to establish the uniform convergence of $F_{\hat{Y}}(y|\mathbf{x})$, which in turn guarantees the uniform convergence of $\hat{r}_{1}(l)$ over the loan decision space.
The Uniform Convergence of $F_{\hat{Y}}(y|\mathbf{x})$
We begin by recalling their assumptions and results.

### v3

@@ 2.2 Functional Risk Formulation
Building on the operational description of the 3PL-led CBEC-SCF model, we now formulate the problem from the perspective of the financial institution. Consider a small- and medium-sized seller engaged in cross-border e-commerce that sells products via an online platform and requires 3PL-led supply chain financing. In this setting, the goods are managed by the 3PL provider during transit, and the revenue generated from eventual sales is primarily allocated to repay the loan. As the financial institution partnering with the 3PL, our goal is to determine an appropriate loan level $l$ based on both the collateral value and projected sales performance, ensuring adequate risk control while aligning financing with the seller’s revenue potential.
@@
where $(x)^{+}=\max\{x,0\}$. Intuitively, $L(l)$ captures the shortfall incurred when the loan level $l$ exceeds the total revenue generated from sales, thereby encapsulating the risk associated with insufficient sales performance. This formulation abstracts the problem to its essential components, enabling us to evaluate how the loss (and hence the risk) varies as a function of the loan level.
The financial institution seeks to evaluate risk measures across the loan decision space $l\in[0,\bar{l}]$, where $\bar{l}$ denotes the maximum loan amount (often equal to the total collateral value to prevent arbitrage opportunities). We introduce a family of key risk measures that vary with $l$, offering flexibility in capturing different dimensions of credit risk. The first is the default probability, which is defined as:
@@
representing the likelihood that sales revenue fails to cover the loan. The second is the expected loss, which is given by:
@@
$\text{LGD}=\frac{\mathbb{E}[L(l)\,|\,L(l)>0]}{l}=\frac{\mathbb{E}[L(l)]}{l\,\Pr\{L(l)>0\}}=\frac{r_{2}(l)}{l\,r_{1}(l)}.$
Moreover, we might be interested in the probability and the expected loss when the loss exceeds a threshold level $\xi$, where $0<\xi<l$. Then,
$\Pr\{L(l)>\xi\}$ $=\Pr\{(l-rY)^{+}>\xi\}$
$=\Pr\{l-rY>\xi\}$
$=\Pr\{Y<(l-\xi)/r\}$
$=r_{1}(l-\xi),$
and
$\mathbb{E}\left[L(l)\,\mathbb{I}\{L(l)>\xi\}\right]=$ $~\mathbb{E}\left[(l-rY)\cdot\mathbb{I}\{l-rY>\xi\}\right]$
$=$ $~\mathbb{E}\left[(l-\xi-rY)\cdot\mathbb{I}\{rY<l-\xi\}\right]$
@@
This demonstrates that $r_{1}(l)$ and $r_{2}(l)$ serve as building blocks for broader risk measures.
In addition, to further broaden the scope of our framework, we propose a more generalized risk measure, $r_{3}(l)$, defined as:
@@
where $g_{l}(\cdot)$ is a loan-level-dependent function that quantifies specific risk forms, and $a(l)$ is a threshold that varies with $l$, often tied to the default point (e.g., $a(l)=l/r$). This formulation further enhances the modeling flexibility, enabling $r_{3}(l)$ to adapt to a wide range of risk measures by tailoring $g_{l}(\cdot)$ and $a(l)$.
  • Recovering basic measures. Both $r_{1}(l)$ and $r_{2}(l)$ are special cases of $r_{3}(l)$. By setting $g_{l}(Y)=1$ and $a(l)=l/r$, we obtain $r_{1}(l)$; and by setting $g_{l}(Y)=l-rY$ and $a(l)=l/r$, we arrive at $r_{2}(l)$.
  • Nonlinear extensions. $r_{3}(l)$ extends to more complex risk measures where the loss function takes on nonlinear forms. For instance, choosing $g_{l}(Y)=(l-rY)^{2}$ and $a(l)=l/r$, we have $r_{3}(l)=\mathbb{E}\left[\left(\left(l-rY\right)^{+}\right)^{2}\right]$, which amplifies larger losses quadratically and emphasizes the severity of extreme shortfalls.
  • Distortion-based measures. Distortion risk measures often take the form of $\mathbb{E}_{\mathbb{Q}}\!\left[(l-rY)^{+}\right]$, where $\mathbb{Q}$ is a new distorted distribution of the sale $Y$ (wang2000class, prelec1998probability). Let $\mathbb{P}$ denote the baseline distribution. Then
    $\mathbb{E}_{\mathbb{Q}}\!\left[(l-rY)^{+}\right]=\mathbb{E}_{\mathbb{P}}\big[\left(l-rY\right)\cdot L(Y)\cdot\mathbb{I}\left\{Y<l/r\right\}\big],$
    where $L(Y)$ denotes the likelihood ratio. In cases where $L(Y)$ may be computed (at least empirically), the distorting risk measure fits into the framework of $r_{3}(l)$ by setting $g_{l}(Y)=(l-rY)\cdot L(Y)$ and $a(l)=l/r$.
  • Revenue-adjusted losses. So far, we assume that the total revenue of the seller is $rY$, which is linear in the sales $Y$. However, sellers may adjust the price of the product based on sales. In that case, we may let $v(Y)$ be the revenue, and the loss becomes lv(Y)l-v(Y). Furthermore, assuming $v(Y)$ is an increasing function of the sales $Y$, then $\left\{l-v(Y)>0\right\}$ is equivalent to $\left\{Y<v^{-1}(l)\right\}$. Then, to compute the expected loss, we may set $g_{l}(Y)=l-v(Y)$ and $a(l)=v^{-1}(l)$ in the definition of $r_{3}(l)$.
Thus, $r_{3}(l)$ provides a unified and versatile framework for capturing sophisticated risk measures, ensuring broad applicability in SCF applications.
Remark 1.
Here, our attention is focused on risk measures expressed in terms of the loan level $l$ and the sales outcome $Y$, which helps clarify the unified formulation of $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$. In the next section, the scope will be broadened to incorporate covariates into the generative model for sales $Y(\mathbf{x})$, where $\mathbf{x}$ may represent product attributes (e.g., ratings, pricing, category), seller characteristics (e.g., review scores, store tenure, fulfillment efficiency), and market factors (e.g., industry trends, buyer preferences, seasonality). Incorporating these features allows us to capture product- and seller-specific heterogeneity, account for broader market influences, and support more refined credit risk assessments, including within-industry comparisons.
Now, the key challenge becomes efficiently computing these risk measures across all $l\in[0,\bar{l}]$. The crux of this challenge is to learn the distribution of $Y$. Suppose we have a generative model $\hat{Y}$ capable of producing $K$ independent and identically distributed observations $\{\hat{Y}_{i}\}_{i=1}^{K}$ rapidly, such that the distribution of $\hat{Y}$ closely approximates that of $Y$. Then, we can estimate these risk measures using Monte Carlo simulation (hong2014monte):
@@
Our proposed functional risk assessment framework revolutionizes traditional SCF credit risk estimation, which typically rely on methods like logistic regression or support vector machines to estimate a single risk measure, often the probability of default, for a fixed loan amount. By modeling risk measures such as $r_{1}(l)$, $r_{2}(l)$, and the versatile $r_{3}(l)$ as functions of loan size $l$, our functional approach provides a unified and comprehensive risk profile across the range of loan levels. Supported by a fast and reliable generative model and the Monte Carlo method, this framework enables financial institutions to estimate diverse credit risk measures and make informed loan sizing decisions, aligning financing with sales potential and significantly enhancing SCF risk management compared to static, single-point traditional methods.
@@ 3 Constructing a Conditional Generative Model for Sales Distribution
In this section, we first formulate the problem of constructing a conditional generative model for sale distribution. We then present QRGMM as an effective generative modeling technique that meets the specific requirements of SCF risk management.
@@ 3.1 Problem Formulation
When a small- and medium-sized seller raises financing through the 3PL-led CEBC-SCF model, the sales revenue of its product will generally be prioritized to repay the loan. Hence, the random variable representing product sales, $Y(\mathbf{x})$, directly determines the loan repayment outcome. Here, the multi-dimensional vector $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{p}$ denotes the associated covariates, such as product features, seller characteristics, and other auxiliary information. We collect $n$ historical observations $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ from e-commerce platforms, where $y_{i}$ are realizations of $Y(\mathbf{x}_{i})$. To effectively manage credit risk in SCF, our objective is to develop a generative model $\hat{Y}(\mathbf{x})$ utilizing the dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. This model should accurately and robustly generate random observations that replicate the conditional distribution of $Y(\mathbf{x})$ given the covariates $\mathbf{x}$, while also ensuring rapid generation.
While generative models have significantly propelled advancements in areas such as image and video synthesis as well as language generation (goodfellow2014generative, radford2018improving, ho2020denoising, liu2024sora), many of them are designed for high-dimensional data, focusing on visual or semantic realism rather than accurately capturing low-dimensional distributions, especially in the tails. Some models like GANs and variational autoencoders (VAE, kingma2013auto) also suffer from mode collapse issues (saxena2021generative), which reduces data diversity and fails to replicate the true distribution accurately, while others (e.g., diffusion-based approaches (ho2020denoising, song2020denoising)) are too slow for large-scale Monte Carlo simulation. In contrast, QRGMM (hong2023learning) is designed to address these issues, offering a compelling choice for SCF applications.
@@ 3.2 The QRGMM Algorithm
  • For $k=1,\ldots,K$, output $\hat{Y}_{k}(\mathbf{x})$ as follows: for $u_{k}\in[\tau_{j},\tau_{j+1})$ with $j=1,\ldots,m{-}2$,
@@
  • Accuracy and Robustness: By explicitly modeling conditional quantiles at multiple levels, QRGMM learns the full shape of a distribution rather than focusing merely on mean predictions or specific performance measures. This quantile-based approach inherently avoids issues such as mode collapse, and it provides more reliable estimates of the tails of the distribution which have more impact on the SCF credit risk. Unlike adversarial models such as GANs that require solving challenging saddle-point optimization (notoriously difficult to tune), QRGMM merely involves fitting quantile regressions and selecting the main parameter, the number of grid points $m$, resulting in a simpler and more stable training process.
@@
  • Speed: Numerical results in hong2023learning show that QRGMM can generate a lot of observations in an extremely short time, e.g., 100,000 random observations in under 0.01 seconds. This efficiency stems from its reliance on simple linear interpolation in the online generation stage, after the quantile regressions have been fitted offline. Such rapid sampling is particularly valuable in SCF risk management, as it enables real-time generation of extensive observations, significantly reducing the computational cost of sample generation. This allows us to generate a substantial number of observations to support the Monte Carlo method, computing various risk measures of interest with enhanced flexibility, while minimizing estimation errors caused by sampling variance.
@@
By framing SCF risk management through the lens of the full conditional distribution of product sales, we underscore the need for a generative model that is not only accurate and robust but also computationally efficient. QRGMM effectively meets these requirements, providing a promising tool for the integrated framework that links credit risk estimation with loan sizing decisions.
However, to fully realize the practical potential of QRGMM in SCF applications, it is essential to go beyond its generative capability and rigorously establish how its learning outcomes translate into reliable credit risk measures estimation. Specifically, building on the problem formulation in Section 2.2, the next critical step, which we address in the following section, is to provide strong theoretical guarantees on the estimators of risk measures. This ensures that the QRGMM-based framework is not only practically effective but also theoretically sound, setting the stage for robust real-world deployment.
Remark 2.
While QRGMM is well suited for univariate sales modeling in the SCF setting, two practical limitations are worth noting for potential future extensions and other applications. First, extending the method to multivariate outcomes is nontrivial because a standard and widely accepted definition of a multivariate quantile function is lacking; sequentially modeling each dimension is possible but may introduce error propagation and slower computation, especially when the dimensionality is high. Second, because QRGMM estimates conditional quantiles on a grid of quantile levels and constructs the full quantile function via interpolation without explicitly imposing monotonicity constraints, the resulting curve may occasionally violate its natural ordering (e.g., a higher quantile taking a smaller estimated value). When such non-monotonic patterns arise, the rearrangement procedure of chernozhukov2010quantile provides a simple post-processing step that restores monotonicity without distorting its statistical validity. A more detailed discussion of these issues of QRGMM can be found in hong2023learning.
4 Asymptotic Analysis of Risk Measures Estimation via QRGMM
@@
In this section, we establish theoretical guarantees for the estimators of $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$ for all $l\in[0,\bar{l}]$, thereby ensuring asymptotic validity for a broad range of SCF risk measures estimation through QRGMM. Prior to proceeding, we emphasize several key considerations.
@@
In this section, we adopt the linear quantile regression model assumption of hong2023learning, which lays the theoretical groundwork for analyzing the asymptotic properties of QRGMM. Suppose $Y(\mathbf{x})$ is a continuous random variable on its support given the input $\mathbf{x}$. Its conditional distribution and density function are denoted as $F_{Y}(\cdot\,|\,\mathbf{x})$ and $f_{Y}(\cdot\,|\,\mathbf{x})$. The linear quantile regression model assumption is stated as follows.
@@
$F^{-1}_{Y}(\tau\,|\,\mathbf{x})=\boldsymbol{\beta}(\tau)^{\intercal}\mathbf{x},~\forall~\tau\in(0,1)$ and $\boldsymbol{\beta}(\tau)$ is continuous in $\tau$.
Assumption 1 posits that the linear quantile regression model accurately represents the true model of the conditional quantile function of $Y$ given $\mathbf{x}$ for all $\tau$ in $(0,1)$, implying that $F^{-1}_{Y}(\tau\,|\,\mathbf{x})$ is linear with respect to the covariates $\mathbf{x}$. Although Assumption 1 may seem restrictive, it is widely used in practice due to its simplicity and interpretability. This assumption is foundational not only because it simplifies the theoretical framework but also because it is underpinned by a substantial literature on classical quantile regression, which provides the necessary theoretical foundation for establishing the asymptotic analysis. Furthermore, in conjunction with domain knowledge, it can be expanded to capture nonlinearity by substituting $\mathbf{x}$ with a set of nonlinear basis functions $\boldsymbol{b(\mathbf{x})}$. In terms of theoretical analysis, the theoretical results can be readily extended when employing basis functions. Supporting evidence from numerical experiments by hong2023learning demonstrates that even when the linear model assumption does not strictly hold, the approximation using basis functions still produces favorable outcomes. In Section 5, we will explore an extension of the linear model to the DeepFM model, enhancing its ability to capture complex combinatorial features.
@@
Sources of Error
Our credit risk measures estimation procedure builds upon QRGMM to produce a generative model $\hat{Y}(\mathbf{x})$ that approximates the distribution of $Y(\mathbf{x})$. Given $\hat{Y}(\mathbf{x})$, we then rely on Monte Carlo simulation to estimate risk measures by generating $K$ observations $\{\hat{Y}_{k}(\mathbf{x})\}_{k=1}^{K}$ from the generative model and transforming them into the estimation of risk measures through (2.2). This process involves three distinct sources of error:
@@
    Interpolation Error: To approximate $F^{-1}_{Y}(\tau\,|\,\mathbf{x})$ for all $\tau$, we rely on a set of $m$ quantile levels and piecewise linear interpolation. This discretization, governed by $m$, introduces interpolation error that diminishes as we increase the number of quantile levels.
@@
    Monte Carlo Sampling Error: After constructing $\hat{Y}(\mathbf{x})$, the final estimation of risk measures is computed by sampling from this approximate distribution. The finite number $K$ of Monte Carlo observations leads to sampling variability, which decreases as $K$ grows.
@@
4.1 Uniform Convergence of Distribution
The asymptotic analysis is based on the uniform convergence of the estimated conditional distribution function $F_{\hat{Y}}(y\,|\,\mathbf{x})$, which serves as the key ingredient for analyzing all subsequent risk measures.
We begin by recalling the assumptions and results of hong2023learning to obtain the uniform convergence of $F_{\hat{Y}}(y\,|\,\mathbf{x})$.